# Trial by FIRE: Probing the dark matter density profile of dwarf galaxies with GraphNPE

### v1

@@ 4.2.1 Result
GravSphere $0.09\pm 0.02$ $0.02\pm 0.01$ $-0.78\pm 0.24$
GraphNPE $0.08\pm 0.02$ $0.01\pm 0.00$ $-0.90\pm 0.14$
$\log_{10}(M_{\mathrm{200m}}$ / $\mathrm{M_{\odot}}$)
@@
The inner profile appears to be more tightly constrained by GravSphere. However, we note that this is likely due to differences in the prior distribution on the inner slope $\gamma$, with $\gamma^{\mathrm{GS}}_{\mathrm{in}}\in[0,1]$ and $\gamma^{\mathrm{gNFW}}\in[-1,2]$. The upper-95% confidence intervals of the density profiles between the two methods tend to be in agreement, though the profiles from GravSphere are slightly steeper in some cases. For the lower-95%, GraphNPE allows the density to flatten more significantly and even decline as $r$ decreases, thus favoring more core-like structure. The decline in the central density is unphysical; however, by allowing $\gamma$ to go negative, GraphNPE avoids running into the prior edge for profiles with pronounced cores.
@@
As in Section 4.1.2, we compute the same mass and structural parameters, i.e. $V_{\mathrm{max}}$, $M_{r_{\star,1/2}}$, and $M_{\mathrm{200m}}^{\mathrm{peak}}$, for each galaxy, with the results shown in Figure 6. Compared to Simple Jeans, the performance of GravSphere closely aligns with that of GraphNPE. Notably, despite differences in modeling choices and prior distributions, the predicted values of $V_{\mathrm{max}}$, $M_{r_{\star,1/2}}$, and $M_{\mathrm{200m}}^{\mathrm{peak}}$ from GravSphere and GraphNPE are remarkably consistent, with their 68% confidence intervals largely overlapping. However, we note that GravSphere consistently underestimates $M_{\mathrm{200m}}^{\mathrm{peak}}$, which is somewhat expected given its preference of steeper outer slopes (Figure 4). Lastly, we evaluate the AE, SE, and NLL performance metrics and present the mean and standard error in Table 4. The performance metrics are comparable for GraphNPE and GravSphere across all mass and structural parameters, with GraphNPE slightly outperforming GravSphere.
@@ 5 Tidal Effects
Figure 8: Recovery of the peak virial mass $M_{\mathrm{200m}}$ of dwarf galaxies in the FIRE-2 dataset. Panels are the same as Figure 7.
@@ 5.1 Quantifying tidal effects with orbital parameters
To quantify the degree of tidal effects, we track the orbital history of each galaxy over the simulation snapshots and compute the following parameters: (1) the number of pericentric passages $N_{\mathrm{peri}}$, (2) the last pericentric distance $r_{\mathrm{peri}}$, (3) the time since the last pericentric passage $t_{\mathrm{peri}}$, and (4) the current distance to host $r_{\mathrm{curr}}$. Satellites that are more tidally disrupted are expected to have lower pericentric distances $r_{\mathrm{peri}}$ and to have completed more orbits around their host galaxies (higher $N_{\mathrm{peri}}$). The latter condition can result from either a high apocentric distance or an early infall time; however, we do not examine these factors separately. The time since the last pericentric passage $t_{\mathrm{peri}}$ provides insight into how long the satellite has had to relax or recover from its most recent tidal interaction, which could influence the persistence of tidal features. Lastly, the current distance $r_{\mathrm{curr}}$ indirectly measures the ongoing tidal forces and also a directly observable quantity. For, a detailed exploration of the orbital dynamics and histories of satellites in Milky Way-mass halos within FIRE-2 simulations, we refer readers to Santistevan et al. (2023).
@@ 5.2 Example: result with $N_{\mathrm{peri}}$
$0.50.2\leq r_{\mathrm{peri}}/R_{\mathrm{host,200m}}^{\mathrm{peri}}<$ 0.5 $63.5\pm 5.4\%$ $94.8\pm 2.6\%$ $72.9\pm 4.9\%$ $96.2\pm 2.2\%$
@@
$10.5\leq r_{\mathrm{curr}}/R_{\mathrm{host,200m}}<$ 1 $67.2\pm 5.2\%$ $93.2\pm 2.8\%$ $71.6\pm 4.9\%$ $96.4\pm 2.0\%$
@@
$52\leq t_{\mathrm{peri}}/\mathrm{Gyr}<$ 5 $61.2\pm 6.2\%$ $93.3\pm 3.3\%$ $77.7\pm 5.4\%$ $91.4\pm 3.8\%$
@@
Figures 7 and 8 present the predicted versus true values (top rows) and residuals (bottom rows) of $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$, respectively, across three different bins of the number of pericentric passages $N_{\mathrm{peri}}=0$, $1$, and $\geq 2$ (left to right columns). Similar figures for other orbital parameters, including $r_{\mathrm{peri}}$, $t_{\mathrm{peri}}$, and $r_{\mathrm{curr}}$, are provided in Appendix C.2. As before, the predicted values and their error bars are the median and 68% confidence intervals of the posterior distributions. Additionally, we distinguish CDM and SIDM galaxies with purple circles and green squares, respectively, with outliers denoted in black.
Figures 7 and 8 clearly illustrate the impact of tides on the predictions of the present-day $V_{\mathrm{max}}$ and the peak $M_{\mathrm{200m}}^{\mathrm{peak}}$. As the number of pericentric passages increase, tidal effects accumulate, and GraphNPE tends to overestimate the $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$. Despite this, we find a good alignment between the predicted and true values in all three bins. In particular, we find that GraphNPE can recover $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within the 68% confidence interval for approximately 64% and 71% of galaxies in the full samples, respectively. Although not shown in the figures, GraphNPE captures approximately 93 % and 95% of galaxies within the 95% confidence interval for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, respectively. Interestingly, this is similar to Genina et al. (2020), which showed that GravSphere recover the enclosed mass distributions to within 68% and 95% confidence for 60% and 90% of their samples, respectively.
@@
As $N_{\mathrm{peri}}$ increases, the coverage fraction decreases, which is expected due to GraphNPE’s tendency to overestimate $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ for more tidally disrupted galaxies. The recovery is overall quite strong: even for galaxies with multiple pericentric passages, we recover their true $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ values within the 68% confidence interval for $57.2\pm 5.1\%$ and $71.2\pm 5.0\%$ of the galaxies, and within the 95% confidence interval for $89.6\pm 3.4\%$ and $93.0\pm 2.8\%$, respectively. We note that the first bin has a lower coverage fraction, which we attribute to a few outlier data points, highlighted as black data points in Figures 7 and 8. These outliers will be discussed in more detail below; here, we note that removing them increases the 68% and 95% coverage fractions in the first $N_{\mathrm{peri}}$ bin to $69.6\%$ and $92.0\%$ for $V_{\mathrm{max}}$ and to $71.6\%$ and $91.9\%$ for $M_{\mathrm{200m}}^{\mathrm{peak}}$.
As noted, a few outlier galaxies in the $N_{\mathrm{peri}}=0$ bin exhibit significantly underestimated $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ values by GraphNPE. These galaxies, highlighted as black data points in Figures 7 and 8, are located at the high-mass end, with true values of $V_{\mathrm{max}}\gtrsim 28,\mathrm{km/s}$ and $\log_{10}M_{\mathrm{200m}}^{\mathrm{peak}}\gtrsim 9.5$, making them less likely to be strongly affected by tidal effects. Upon closer examination, we find that these galaxies are all nearing their first pericentric passage and are likely experiencing significant tidal shocks due to their proximity to the host galaxy. This example underscores a key limitation of $N_{\mathrm{peri}}$: while it effectively tracks cumulative tidal effects, it does not capture instances where a galaxy is undergoing strong, transient tidal interactions. Therefore, to gain a more comprehensive understanding of the model’s performance and the complexity of tidal interactions, it is crucial to consider additional orbital parameters, as discussed in Section 5.3.
@@
To briefly summarize, we emphasize that, despite these challenges, GraphNPE’s overall performance remains strong. In particular, GraphNPE recovers $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within the 68% and 95% confidence intervals for a large fraction of galaxies. This accuracy matches that of the previous version of GravSphere in Genina et al. (2020) for enclosed mass distributions, while here we specifically demonstrate similar reliability for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$. The strong performance on $M_{\mathrm{200m}}^{\mathrm{peak}}$ suggests that (1) present-day stellar kinematics encode sufficient information to infer the peak halo mass and (2) GraphNPE is capable of effectively extracting this information. The former is somewhat unsurprising, as past studies have shown that $M_{\mathrm{200m}}^{\mathrm{peak}}$ can be recovered if the tidal radius is sufficient far from the stellar half-stellar mass radius of satellites (e.g. Read et al., 2006; Errani et al., 2018; Read et al., 2018; Read & Erkal, 2019). This result is further highlighted by the fact that the simple Jeans model in Section 4.1 appears to overestimate the peak mass more consistently, as shown from the samples in Figure 3.
@@ 5.3 Result with all orbital parameters
We now examine the performance of GraphNPE for the remaining orbital parameters: namely, the last pericentric distance $r_{\mathrm{peri}}$, the current distance $r_{\mathrm{curr}}$, and the time since last pericenter $t_{\mathrm{peri}}$. As before, Table 5 presents the 68% and 95% coverage fractions, now calculated across three separate bins for each of the above parameters, resulting in a total of nine bins.
@@
Interestingly, GraphNPE appears to be conservative in estimating $M_{\mathrm{200m}}^{\mathrm{peak}}$, with overly large confidence intervals that cover more samples than expected. In some cases (e.g., $t_{\mathrm{peri}}>5\,\mathrm{Gyr}$, $N_{\mathrm{peri}}=1$), GraphNPE can recover $M_{\mathrm{200m}}^{\mathrm{peak}}$ for close to 100% of the populations. The broad confidence intervals might arise from extrapolating the DM density profiles when computing $M_{\mathrm{200m}}^{\mathrm{peak}}$, which introduce additional uncertainties. This issue could be mitigated with a more accurate model for the profile tails or a more extended tracer population.
@@
To better quantify the accuracy of GraphNPE on predicting $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, we compute their mean residuals across different bins of the four orbital parameters. We first note an important consideration. It is evident from Figures 7 and 8 that there are distinct populations of galaxies within the three $N_{\mathrm{peri}}$ bins. Specifically, more massive galaxies–both in their peak and present-day masses–are more likely to occupy lower $N_{\mathrm{peri}}$ bin. Galaxies with high $N_{\mathrm{peri}}$ likely fell into their host halos at earlier times, resulting in generally lower values of $M_{\mathrm{200m}}^{\mathrm{peak}}$, as they had less time to accrete mass before being incorporated into their host. Additionally, these galaxies are expected to have experienced more stripping, leading to lower values of present-day $V_{\mathrm{max}}$. It is therefore crucial to disentangle tidal effects from those of galaxy populations. More massive galaxies tend to host more stars, and the tighter constraints by GraphNPE observed in the lower $N_{\mathrm{peri}}$ bin could arise either from the absence of significant tidal effects or from the larger number of stars. Therefore, for each bin of the tidal parameter, we further subdivide the galaxies into three equal-sized bins based on $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$. Each sub-bin contains at least 10 galaxies to ensure statistical robustness in our analysis.
@@
Figures 9 and 10 present the fractional residuals for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, respectively, across different bins of $N_{\mathrm{peri}}$, $r_{\mathrm{peri}}$, $r_{\mathrm{curr}}$, and $t_{\mathrm{peri}}$ (left to right columns). The top row shows the net (signed) residuals, while the bottom row shows the absolute (unsigned) residuals. In each panel, the color and marker indicate the bins of the corresponding tidal parameter, with blue circles, yellow squares, and green triangles representing bins in order from least to most likely to be tidally disrupted. The $y$-position of each data point represents the mean residuals of galaxies in the bin, with the vertical error bars showing the standard errors in estimating the mean. The $x$-position corresponds to the mean of the true $V_{\mathrm{max}}$ or $M_{\mathrm{200m}}^{\mathrm{peak}}$ values for all galaxies in the bin, while the horizontal error bars indicate the range of true $V_{\mathrm{max}}$ or $M_{\mathrm{200m}}^{\mathrm{peak}}$ values within the bin. We do not include the outliers discussed previously in the calculation.
@@
The fractional residual of $V_{\mathrm{max}}$ exhibits similar trends across the panels of Figure 9. Within the same tidal bin, GraphNPE overestimates $V_{\mathrm{max}}$ for galaxies at low $V_{\mathrm{max}}$, as expected since these galaxies are more susceptible to tides. As $V_{\mathrm{max}}$ increases, both the net residual and absolute residual steadily decrease. Beyond $V_{\mathrm{max}}\gtrsim 20\,\mathrm{km/s}$, the net residual approaches zero, while the absolute residual flattens out at approximately $10-20\%$. When comparing between different tidal bins, we find that galaxies in the more disrupted bins tend to have lower values of $V_{\mathrm{max}}$, consistent with the findings in Figure 7. At similar $V_{\mathrm{max}}$ values, the mean residuals tend to be larger in more disrupted bins, although there is significant scatter in both $V_{\mathrm{max}}$ and residuals within each bin. For $V_{\mathrm{max}}$ in the range $10-20\,\mathrm{km/s}$, the absolute residuals are typically around $20\%$, peaking at approximately $25\pm 10\%$ for galaxies with $V_{\mathrm{max}}\approx 11\,\mathrm{km/s}$ in the $r_{\mathrm{peri}}<0.2\,R_{\mathrm{host,200m}}^{\mathrm{peri}}$ bin. At $V_{\mathrm{max}}\gtrsim 20\,\mathrm{km/s}$, both the net and absolute residuals flatten across all tidal bins, as also previously noted, resulting in no significant performance differences.

### v2

@@ 4.2.1 Result
GravSphere $09\pm 0.02 0.0$$02\pm 0.01$ $0.78\pm 0.24$
GraphNPE $08\pm 0.02 0.0$$01\pm 0.00$ $0.90\pm 0.14$
$\log_{10}(M_{\mathrm{200m}}^{\mathrm{peak}}$ / $\mathrm{M_{\odot}}$)
@@
The inner profile appears to be more tightly constrained by GravSphere. However, we note that this is likely due to differences in the prior distribution on the inner slope $\gamma$, with $\gamma^{\mathrm{GS}}_{\mathrm{in}}\in[0,1]$ and $\gamma^{\mathrm{gNFW}}\in[-1,2]$. The upper-95% confidence intervals of the density profiles between the two methods tend to be in agreement, though the profiles from GravSphere are slightly steeper in some cases. For the lower-95%, GraphNPE allows the density to flatten more significantly and even decline as $r$ decreases, thus favoring more core-like structures. The decline in the central density is unphysical; however, by allowing $\gamma$ to go negative, GraphNPE avoids running into the prior edge for profiles with pronounced cores.
@@
As in Section 4.1.2, we compute the same mass and structural parameters, i.e. $V_{\mathrm{max}}$, $M_{r_{\star,1/2}}$, and $M_{\mathrm{200m}}^{\mathrm{peak}}$, for each galaxy, with the results shown in Figure 7. Compared to Simple Jeans, the performance of GravSphere closely aligns with that of GraphNPE. Notably, despite differences in modeling choices and prior distributions, the predicted values of $V_{\mathrm{max}}$, $M_{r_{\star,1/2}}$, and $M_{\mathrm{200m}}^{\mathrm{peak}}$ from GravSphere and GraphNPE are remarkably consistent, with their 68% confidence intervals largely overlapping. However, we note that GravSphere consistently underestimates $M_{\mathrm{200m}}^{\mathrm{peak}}$, which is somewhat expected given its preference for steeper outer slopes (Figure 5). Additionally, we evaluate the AE, SE, and NLL performance metrics and present the mean and standard error in Table 4. The performance metrics are comparable for GraphNPE and GravSphere across all mass and structural parameters, with GraphNPE slightly outperforming GravSphere.
Lastly, Figure 8 shows the fractional residuals of the mass and structural parameters as a function of $N_{\star}$. Unlike the comparison with Simple Jeans, it is more difficult to draw definitive conclusions about GraphNPE versus GravSphere performance due to the limited sample size. However, the results suggest that GraphNPE may also provide a modest advantage over GravSphere at low tracer counts, at least in the case of $V_{\mathrm{max}}$.
@@ 5 Tidal Effects
Figure 10: Recovery of the peak virial mass $M_{\mathrm{200m}}^{\mathrm{peak}}$ of dwarf galaxies in the FIRE-2 dataset. Panels are the same as Figure 9.
@@ 5.1 Quantifying tidal effects with orbital parameters
To quantify the degree of tidal effects, we track the orbital history of each galaxy over the simulation snapshots and compute the following parameters: (1) the number of pericentric passages $N_{\mathrm{peri}}$, (2) the last pericentric distance $r_{\mathrm{peri}}$, (3) the time since the last pericentric passage $t_{\mathrm{peri}}$, and (4) the current distance to host $r_{\mathrm{curr}}$. Satellites that are more tidally disrupted are expected to have lower pericentric distances $r_{\mathrm{peri}}$ and to have completed more orbits around their host galaxies (higher $N_{\mathrm{peri}}$). The latter condition can result from either a shorter orbital period or an early infall time; however, we do not examine these factors separately. The time since the last pericentric passage $t_{\mathrm{peri}}$ provides insight into how long the satellite has had to relax or recover from its most recent tidal interaction, which could influence the persistence of tidal features. Lastly, the current distance $r_{\mathrm{curr}}$ indirectly measures the ongoing tidal forces and also a directly observable quantity. For, a detailed exploration of the orbital dynamics and histories of satellites in Milky Way-mass halos within FIRE-2 simulations, we refer readers to Santistevan et al. (2023).
@@ 5.2 Example: result with $N_{\mathrm{peri}}$
$0.2\leq r_{\mathrm{peri}}/R_{\mathrm{host,200m}}^{\mathrm{peri}}<0.5$ $63.5\pm 5.4\%$ $94.8\pm 2.6\%$ $72.9\pm 4.9\%$ $96.2\pm 2.2\%$
@@
$0.5\leq r_{\mathrm{curr}}/R_{\mathrm{host,200m}}<1$ $67.2\pm 5.2\%$ $93.2\pm 2.8\%$ $71.6\pm 4.9\%$ $96.4\pm 2.0\%$
@@
$2\leq t_{\mathrm{peri}}/\mathrm{Gyr}<5$ $61.2\pm 6.2\%$ $93.3\pm 3.3\%$ $77.7\pm 5.4\%$ $91.4\pm 3.8\%$
@@
Figures 9 and 10 present the predicted versus true values (top rows) and residuals (bottom rows) of $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}$, respectively, across three different bins of the number of pericentric passages $N_{\mathrm{peri}}=0$, $1$, and $\geq 2$ (left to right columns). Similar figures for other orbital parameters, including $r_{\mathrm{peri}}$, $t_{\mathrm{peri}}$, and $r_{\mathrm{curr}}$, are provided in Appendix C.3. As before, the predicted values and their error bars are the median and 68% confidence intervals of the posterior distributions. Additionally, we distinguish CDM and SIDM galaxies with purple circles and green squares, respectively, with outliers denoted in black.
Figures 9 and 10 clearly illustrate the impact of tides on the predictions of the present-day $V_{\mathrm{max}}$ and the peak $M_{\mathrm{200m}}^{\mathrm{peak}}$. As the number of pericentric passages increase, tidal effects accumulate, and GraphNPE tends to overestimate the $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$. Despite this, we find a good alignment between the predicted and true values in all three bins. In particular, we find that GraphNPE can recover $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within the 68% confidence interval for approximately 64% and 71% of galaxies in the full samples, respectively. Although not shown in the figures, GraphNPE captures approximately 93 % and 95% of galaxies within the 95% confidence interval for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, respectively. Interestingly, this is similar to Genina et al. (2020), which showed that GravSphere recover the enclosed mass distributions to within 68% and 95% confidence for 60% and 90% of their samples, respectively.
@@
As $N_{\mathrm{peri}}$ increases, the coverage fraction decreases, which is expected due to GraphNPE’s tendency to overestimate $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ for more tidally disrupted galaxies. The recovery is overall quite strong: even for galaxies with multiple pericentric passages, we recover their true $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ values within the 68% confidence interval for $57.2\pm 5.1\%$ and $71.2\pm 5.0\%$ of the galaxies, and within the 95% confidence interval for $89.6\pm 3.4\%$ and $93.0\pm 2.8\%$, respectively. We note that the first bin has a lower coverage fraction, which we attribute to a few outlier data points, highlighted as black data points in Figures 9 and 10. These outliers will be discussed in more detail below; here, we note that removing them increases the 68% and 95% coverage fractions in the first $N_{\mathrm{peri}}$ bin to $69.6\%$ and $92.0\%$ for $V_{\mathrm{max}}$ and to $71.6\%$ and $91.9\%$ for $M_{\mathrm{200m}}^{\mathrm{peak}}$.
As noted, a few outlier galaxies in the $N_{\mathrm{peri}}=0$ bin exhibit significantly underestimated $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ values by GraphNPE. These galaxies, highlighted as black data points in Figures 9 and 10, are located at the high-mass end, with true values of $V_{\mathrm{max}}\gtrsim 28,\mathrm{km/s}$ and $\log_{10}M_{\mathrm{200m}}^{\mathrm{peak}}\gtrsim 9.5$, making them less likely to be strongly affected by tidal effects. Upon closer examination, we find that these galaxies are all nearing their first pericentric passage and are likely experiencing significant tidal shocks due to their proximity to the host galaxy. This example underscores a key limitation of $N_{\mathrm{peri}}$: while it effectively tracks cumulative tidal effects, it does not capture instances where a galaxy is undergoing strong, transient tidal interactions. Therefore, to gain a more comprehensive understanding of the model’s performance and the complexity of tidal interactions, it is crucial to consider additional orbital parameters, as discussed in Section 5.3.
@@
To briefly summarize, we emphasize that, despite these challenges, GraphNPE’s overall performance remains strong. In particular, GraphNPE recovers $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$ within the 68% and 95% confidence intervals for a large fraction of galaxies. This accuracy matches that of the previous version of GravSphere in Genina et al. (2020) for enclosed mass distributions, while here we specifically demonstrate similar reliability for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$. The strong performance on $M_{\mathrm{200m}}^{\mathrm{peak}}$ suggests that (1) present-day stellar kinematics encode sufficient information to infer the peak halo mass and (2) GraphNPE is capable of effectively extracting this information. The former is somewhat unsurprising, as past studies have shown that $M_{\mathrm{200m}}^{\mathrm{peak}}$ can be recovered if the tidal radius is sufficiently far from the half-stellar mass radius of satellites (e.g. Read et al., 2006; Errani et al., 2018; Read et al., 2018; Read & Erkal, 2019). This result is further highlighted by the fact that the simple Jeans model in Section 4.1 appears to overestimate the peak mass more consistently, as shown from the samples in Figure 3.
@@ 5.3 Result with all orbital parameters
We now examine the performance of GraphNPE for the remaining orbital parameters: namely, the last pericentric distance $r_{\mathrm{peri}}$, the current distance $r_{\mathrm{curr}}$, and the time since the last pericenter $t_{\mathrm{peri}}$. As before, Table 5 presents the 68% and 95% coverage fractions, now calculated across three separate bins for each of the above parameters, resulting in a total of nine bins.
@@
Interestingly, GraphNPE appears to be conservative in estimating $M_{\mathrm{200m}}^{\mathrm{peak}}$, with overly large confidence intervals that cover more samples than expected. In some cases (e.g., $t_{\mathrm{peri}}>5\,\mathrm{Gyr}$, $N_{\mathrm{peri}}=1$), GraphNPE can recover $M_{\mathrm{200m}}^{\mathrm{peak}}$ for close to 100% of the populations. The broad confidence intervals might arise from extrapolating the DM density profiles when computing $M_{\mathrm{200m}}^{\mathrm{peak}}$, which introduces additional uncertainties. This issue could be mitigated with a more accurate model for the profile tails or a more extended tracer population.
@@
To better quantify the accuracy of GraphNPE on predicting $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, we compute their mean residuals across different bins of the four orbital parameters. We first note an important consideration. It is evident from Figures 9 and 10 that there are distinct populations of galaxies within the three $N_{\mathrm{peri}}$ bins. Specifically, more massive galaxies–both in their peak and present-day masses–are more likely to occupy lower $N_{\mathrm{peri}}$ bin. Galaxies with high $N_{\mathrm{peri}}$ likely fell into their host halos at earlier times, resulting in generally lower values of $M_{\mathrm{200m}}^{\mathrm{peak}}$, as they had less time to accrete mass before being incorporated into their host. Additionally, these galaxies are expected to have experienced more stripping, leading to lower values of present-day $V_{\mathrm{max}}$. It is therefore crucial to disentangle tidal effects from those of galaxy populations. More massive galaxies tend to host more stars, and the tighter constraints by GraphNPE observed in the lower $N_{\mathrm{peri}}$ bin could arise either from the absence of significant tidal effects or from the larger number of stars. Therefore, for each bin of the tidal parameter, we further subdivide the galaxies into three equal-sized bins based on $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$. Each sub-bin contains at least 10 galaxies to ensure statistical robustness in our analysis.
@@
Figures 11 and 12 present the fractional residuals for $V_{\mathrm{max}}$ and $M_{\mathrm{200m}}^{\mathrm{peak}}$, respectively, across different bins of $N_{\mathrm{peri}}$, $r_{\mathrm{peri}}$, $r_{\mathrm{curr}}$, and $t_{\mathrm{peri}}$ (left to right columns). The top row shows the net (signed) residuals, while the bottom row shows the absolute (unsigned) residuals. In each panel, the color and marker indicate the bins of the corresponding tidal parameter, with blue circles, yellow squares, and green triangles representing bins in order from least to most likely to be tidally disrupted. The $y$-position of each data point represents the mean residuals of galaxies in the bin, with the vertical error bars showing the standard errors in estimating the mean. The $x$-position corresponds to the mean of the true $V_{\mathrm{max}}$ or $M_{\mathrm{200m}}^{\mathrm{peak}}$ values for all galaxies in the bin, while the horizontal error bars indicate the range of true $V_{\mathrm{max}}$ or $M_{\mathrm{200m}}^{\mathrm{peak}}$ values within the bin. We do not include the outliers discussed previously in the calculation.
@@
The fractional residual of $V_{\mathrm{max}}$ exhibits similar trends across the panels of Figure 11. Within the same tidal bin, GraphNPE overestimates $V_{\mathrm{max}}$ for galaxies at low $V_{\mathrm{max}}$, as expected since these galaxies are more susceptible to tides. As $V_{\mathrm{max}}$ increases, both the net residual and absolute residual steadily decrease. Beyond $V_{\mathrm{max}}\gtrsim 20\,\mathrm{km/s}$, the net residual approaches zero, while the absolute residual flattens out at approximately $10-20\%$. When comparing different tidal bins, we find that galaxies in the more disrupted bins tend to have lower values of $V_{\mathrm{max}}$, consistent with the findings in Figure 9. At similar $V_{\mathrm{max}}$ values, the mean residuals tend to be larger in more disrupted bins, although there is significant scatter in both $V_{\mathrm{max}}$ and residuals within each bin. For $V_{\mathrm{max}}$ in the range $10-20\,\mathrm{km/s}$, the absolute residuals are typically around $20\%$, peaking at approximately $25\pm 10\%$ for galaxies with $V_{\mathrm{max}}\approx 11\,\mathrm{km/s}$ in the $r_{\mathrm{peri}}<0.2\,R_{\mathrm{host,200m}}^{\mathrm{peri}}$ bin. At $V_{\mathrm{max}}\gtrsim 20\,\mathrm{km/s}$, both the net and absolute residuals flatten across all tidal bins, as also previously noted, resulting in no significant performance differences.